OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = –32 and a9 = –120?

OpenStudy (anonymous):

Difference between terms is 11. So, the series is: a(n) = -32-(11*(n-1)) a(32) = -32-11*31 = -373

OpenStudy (anonymous):

thank you so much for your help (:

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

im new to this, so im kind of lost on what that means lol